Please help!!
Can someone help me solve the math problem below please?
4x+3=3x+6
Thank you!!!!

Answer by mananth(16949)

4x+3=3x+6
...
bring the 3x to the left side
add -3x
4x+3-3x= 3x-3x+6
3x -3x cancel off
4x-3x +3 = 6
x+3 = 6
3 to go to other side
add -3 to both sides
x+3-3 = 6-3
+3,-3 cancel off
x= 3
...
m.ananth@hotmail.ca